OpenStudy (anonymous):

Write the expression below so that only a single logarithm or exponential function appears. (1/2)ln(z) − ln(5 + x) − 4 ln(y)

zepdrix (zepdrix):

Rule of logs:\[\Large \color{royalblue}{b\cdot\log(a) \quad=\quad \log(a^b)}\]Understand how we can use this rule to deal with the 1/2 coefficient on the first term?

OpenStudy (anonymous):

ln(z)^(1/2)

zepdrix (zepdrix):

The exponent is being applied to the z, not the log. Just be careful the way you write that :) ln(z^(1/2))

zepdrix (zepdrix):

We can do the same with the 4, yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ln(z^(1/2))-ln(5-x)-ln(y^4)

zepdrix (zepdrix):

\[\Large \ln \left(z^{1/2}\right) − \ln(5 + x) − \ln\left(y^4\right)\]Ok cool.

zepdrix (zepdrix):

We'll want to apply another rule of logs from here:\[\Large \color{#CC0033}{\log(a)-\log(b) \quad=\quad \log\left(\frac{a}{b}\right)}\]

OpenStudy (anonymous):

I get this ln((z^(1/2))/((5+x)(y^4)))

zepdrix (zepdrix):

\[\Large \ln\left[\frac{z^{1/2}}{(5+x)y^4}\right]\] With the last two terms in the denominator? Mmmm yah that looks right. Good job! :)

OpenStudy (anonymous):

yay thank you:)

OpenStudy (anonymous):

hey zep could u help me with another problem.. i dont know what im doing wrong

zepdrix (zepdrix):

sure

OpenStudy (anonymous):

A very small tumor of initial size M is observed every ten days, and over each ten-day period, it has grown 70% larger. (a) Write a formula representing the size of the tumor after t days (t = 0, 10, 20, . . .)

OpenStudy (anonymous):

(b) On which day will we finally observe that the tumor is more than 10 times its original size? (Assume that the observations are strictly periodically conducted as specified.)

OpenStudy (anonymous):

for part a) I'm getting (1.7)^tM and it says is wrong

zepdrix (zepdrix):

Is the M in the exponent? I can't tell with the way you wrote it :3

OpenStudy (anonymous):

no multiplying

OpenStudy (anonymous):

(1.7^t)*M

zepdrix (zepdrix):

Mmm lemme see if I've got the right idea here.\[\Large f(t)=Mb^t\]Where M is the initial mass of the tumor. at \(\Large t=10\) the tumor has grown to a size of 170% or \(\Large 1.7M\) \[\Large 1.7M=Mb^{10}\]

zepdrix (zepdrix):

\[\Large 1.7=b^{10}\]

zepdrix (zepdrix):

\[\Large b\approx 1.054\]

zepdrix (zepdrix):

Do you have a way to check this? To make sure I'm not doing something really stupid :) lol

OpenStudy (anonymous):

well i need a formula for part a

OpenStudy (anonymous):

and i get (1.7^t)M and no i dont have a way to check it :(

zepdrix (zepdrix):

The way I did it, we would end up with,\[\Large b=1.7^{1/10}\]Plugging this into our function gives us,\[\Large f(t)=M\left(1.7^{1/10}\right)^t \qquad\to\qquad f(t)=M\left(1.7\right)^{t/10}\]

zepdrix (zepdrix):

^ Does this way make sense? :o How did you come up with b=1.7?

OpenStudy (anonymous):

yup that worked!

OpenStudy (anonymous):

you are amazing!

zepdrix (zepdrix):

yay team \c:/

OpenStudy (anonymous):

hey for part b do i just plug 10 for t?

OpenStudy (anonymous):

and solve for M?

zepdrix (zepdrix):

On which day, \(\large t\), will we finally observe that the tumor, \(\large f(t)\), is more than 10 times, \(\large \gt 10M\) its original size?

zepdrix (zepdrix):

\[\Large 10M<M\left(1.7\right)^{t/10}\]

zepdrix (zepdrix):

I guess we don't really need the inequality. We can just use an = sign. Whatever t day we end up with, we will round up to the next day to show the day when the mass is 10times GREATER THAN it's initial size.

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

t is about 43.3936

OpenStudy (anonymous):

so 4 days?

OpenStudy (anonymous):

44

zepdrix (zepdrix):

ya that sounds right! :)

OpenStudy (anonymous):

says its wrong :(

zepdrix (zepdrix):

Hmm thinkinggg

zepdrix (zepdrix):

(Assume that the observations are strictly `periodically` conducted as specified.) Only periodically? So does this mean our t MUST BE a multiple of 10 maybe? Maybe try t=50 :\ If not we'll have to think about it a lil more.

OpenStudy (anonymous):

that worked!!!!

OpenStudy (anonymous):

thank you sooo much!

zepdrix (zepdrix):

Oh cool! :) Strange wording on the problem. heh

zepdrix (zepdrix):

\[\Large f(\color{#CC0033}{x})=4-4\color{#CC0033}{x}-\color{#CC0033}{x}^2\] \[\Large f(\color{#CC0033}{3})=4-4\cdot\color{#CC0033}{3}-\color{#CC0033}{3}^2\]\[\Large f(\color{#CC0033}{3+h})=4-4(\color{#CC0033}{3+h})-(\color{#CC0033}{3+h})^2\]

zepdrix (zepdrix):

The next part is a little tricky. You need to feel somewhat comfortable with function notation. We're going to plug in the pieces.\[\Large \frac{f(\color{#CC0033}{3+h})-f(\color{#CC0033}{3})}{h}\]

OpenStudy (anonymous):

yea im having trouble with this part

zepdrix (zepdrix):

Plugging it all in is a bit of a pain in the butt. We get something like this,\[\large \frac{4-4(3+h)-(3+h)^2-\left[4-4(3)-3^2\right]}{h}\]

zepdrix (zepdrix):

Notice how I put parentheses around the f(3)? That's important. The negative needs to be distributed to each term in the brackets.

OpenStudy (anonymous):

ok so now i solve for h

zepdrix (zepdrix):

Here's a little tip to try to keep in mind. Store this in the back of your head somewhere, it might come in handy later. `If you do all of your simplification correctly, every term that doesn't contain an h will cancel out`.

zepdrix (zepdrix):

Not solve for h, we're just trying to simplify the expression.

zepdrix (zepdrix):

Do a bunch of multiplication, expand out the binomial with the square on it, and simplify! :D

zepdrix (zepdrix):

Remember how to expand this out? \(\Large (3+h)^2\)

OpenStudy (anonymous):

omg this is going to be a pain... let me try this lol

OpenStudy (anonymous):

i get (-36-3h+h^2)/h

OpenStudy (anonymous):

not sure on that

zepdrix (zepdrix):

Remember, EVERYTHING that doesn't have an `h` in the numerator should have cancelled out. So you shouldn't be left with -36. Hmm

zepdrix (zepdrix):

If we ignore the denominator for a sec,\[\large 4-4(3+h)-(3+h)^2-\left[4-4(3)-3^2\right]\]So we get some stuff,\[\large 4-12-4h-9-6h-h^2-4+12+9\]

zepdrix (zepdrix):

See any mistakes you may have made? :o

OpenStudy (anonymous):

-12-4h-6h-h^2?

OpenStudy (anonymous):

-12-10h-h^2?

zepdrix (zepdrix):

Where is the -12 coming from? :o See the cancellations?\[\large \cancel4\cancel{-12}-4h\cancel{-9}-6h-h^2\cancel{-4}\cancel{+12}\cancel{+9}\]

OpenStudy (anonymous):

omg im blind

zepdrix (zepdrix):

:3

OpenStudy (anonymous):

(-10h-h^2)/h

OpenStudy (anonymous):

can i simplify more than that?

zepdrix (zepdrix):

yah you can divide an h out of every term.

zepdrix (zepdrix):

Or you can factor an h out of each term in the numerator, if you feel more comfortable doing that step before the division.

OpenStudy (anonymous):

-10-h is the final answer?

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

yay thank you!

OpenStudy (anonymous):

I have a similar question.. could u set it up for me and let me try to do it?

zepdrix (zepdrix):

sure

zepdrix (zepdrix):

working on b?

zepdrix (zepdrix):

I hate that notation for composition, let's use this instead.\[\Large (g\circ f)(4) \quad=\quad g\left[f(4)\right]\]

OpenStudy (anonymous):

shouldnt it be -3?

zepdrix (zepdrix):

Work from the inside out.\[\Large g\left[\color{#F35633}{f(4)}\right]\] \[\Large \color{#F35633}{f(4)=3}\]\[\Large g\left[\color{#F35633}{f(4)}\right] \quad=\quad g\left[\color{#F35633}{3}\right] \quad=\quad ?\]

OpenStudy (anonymous):

1?

zepdrix (zepdrix):

yup!

OpenStudy (anonymous):

omg!!! ty

OpenStudy (anonymous):

u make it look so easy

OpenStudy (anonymous):

could you help me with part e) that looks complicated

zepdrix (zepdrix):

\[\Large h(\color{#3399AA}{x})=\sqrt{\color{#3399AA}{x}^2-f(g(\color{#3399AA}{x}))}\]

zepdrix (zepdrix):

\[\Large h(\color{#3399AA}{2})=\sqrt{\color{#3399AA}{2}^2-f(g(\color{#3399AA}{2}))}\]

zepdrix (zepdrix):

Again, work from the inside out,\[\Large g(2)=?\]

OpenStudy (anonymous):

5?

zepdrix (zepdrix):

\[\Large h(\color{#3399AA}{2})=\sqrt{\color{#3399AA}{2}^2-f(5)}\]

zepdrix (zepdrix):

f(5)=? :o

OpenStudy (anonymous):

f(5)=4

OpenStudy (anonymous):

sqrt(4-4)

OpenStudy (anonymous):

undefined?

zepdrix (zepdrix):

Mmm sqrt0 = 0 cuz 0*0=0

zepdrix (zepdrix):

Or better said: 0^2=0

OpenStudy (anonymous):

it says 0 is not the correct answer

zepdrix (zepdrix):

Hmmm :\